# U2A: Unified Unimodal Adaptation for Robust and Efficient Multimodal Learning

### v1

@@ 3 U2A: Unified Unimodal Adaptation
where $\lambda$ is a hyperparameter. In all our experiments, we set $\lambda=0.20$.
In summary, we use pretrained encoders for each modality and jointly train the modality-specific adaptation parameters ($\Delta_{f,m}$) and the classifier head on the multimodal dataset $\mathcal{D}$. While training our model on complete multimodal data, we apply random modality dropout. Specifically, for each training iteration, we either use both modalities or randomly drop one. This allows the model to leverage the learned mask tokens for making accurate predictions when a modality is missing. We do not apply random modality dropout when the training set has missing modalities, as the model naturally encounters opportunities to use the mask tokens for the samples with missing modalities.
4 Experimental Results
4.1 Datasets
We evaluate our approach on five popular multimodal datasets across different tasks. A brief description of each dataset is provided here, with further details in Section 1 in the supplementary.
4.2 Implementation Details
Audio-Video datasets. We use AST model [57] pretrained on AudioSet dataset [58] as audio encoder and ViT-B [55] from CLIP model [56] as video encoder. Raw audio is converted to a 128-bin mel spectrogram using a 16 kHz sampling rate, with a maximum length of 1024 frames. For video, we sample 3 frames randomly while training and uniformly while testing. Other pre-processing steps are similar to [36, 33, 37].
Hyperparameter settings. We use Python 3.8.19 and PyTorch 2.2.2 for training and evaluating our models. All the models are trained using two NVIDIA RTX 2080Ti GPUs. For vision-language datasets, we set the learning rate to $10^{-3}$ and train the models for 10 epochs with a batch size of 8. For audio-video datasets, the learning rate is set to $5\times 10^{-5}$ and models are trained for 100 epochs with a batch size of 4. We use AdamW [59] optimizer with $\epsilon=10^{-8}$ and weight decay = 0.02. Cross entropy loss is applied while training the models. We utilize a polynomial learning rate scheduler with power=0.9 and set the first 5 epoch as warm-up. The learning rate is set to 0.1 times the original learning rate during warm-up epochs. We set LoRA [27] rank = 1 and insert them after query, key, value and output layers of each transformer block. Further details with all the hyperparameters can be found in Section 2 in the supplementary.
For each task and dataset, we compare against existing works when published results are available. Due to this criterion, the set of baselines varies across experiments based on the availability of published results. In each table, the best and second best results are bold and underlined, respectively.
Table 1: Performance comparison on UPMC Food-101 and MM-IMDb dataset when both modalities are available during training and testing. We report % accuracy for UPMC Food-101 and F1-Macro/F1-Micro score for MM-IMDb dataset. ‘-’ indicates that these values are not reported in the corresponding papers.
Methods Parameters (M) UPMC MM-IMDb
Total Learnable Food-101
ViLT 112.63 112.63 92.00 55.30 / 64.70
MBT 196.00 196.00 93.56 59.60 / 64.81
MMBT 170.00 170.00 94.10 60.80 / 66.10
MMLoRA 196.10 196.10 93.70 61.70 / 67.20
PMF 198.43 2.54 91.51 58.77 / 64.51
PMF-L 643.95 4.44 91.68 61.66 / 66.72
MLA 218.26 218.26 93.33 - / -
Ours 195.51 0.27 94.47 63.58 / 69.23
4.3 Complete Modality Performance
We train and evaluate our method in the modality-complete scenario, where both modalities are available during training and testing. We compare our results with (1) baseline multimodal models: MBT [30], MMBT [28], ViLT [8]; (2) prompting based methods: PMF [60]; (3) modulation and gating based methods: FiLM [61] and Bilinear-Gated (BiGated) [62] (4) model finetuning based approach: MMLoRA [37] and (5) gradient modulation based methods: G-Blend [32], OGM-GE [33], MLA [36].
As summarized in Table 1, our approach outperforms all the baseline methods in both UPMC Food-101 and MM-IMDb datasets. Similarly, our method shows comparable or better performance in all the audio-video datasets as shown in Table 2. MMLoRA [37] shows close performance to our method and also bears some similarities and key differences. MMLoRA first finetunes the unimodal models on unimodal data, followed by a multimodal adaptation on complete multimodal data. In contrast, our approach freezes the encoders pretrained on unimodal data and learns only the low-rank adapters, which reduces the total number of learnable parameters significantly while achieving superior performance in most of the datasets. Furthermore, our method achieves robust performance for missing modalities, as we discuss in the next section.
4.4 Missing Modality Performance
We consider the following two general cases when evaluating our approach in missing modality scenarios.
4.4.1 Missing modality during both training and testing
We follow the experimental setup from [24, 25] and train and test our models with same modality ratio. Let $\eta$ denote the percentage of samples with missing modalities. For bi-modal vision-language tasks, we consider three scenarios: missing-text, missing-image, and missing-both. For the first two cases, $\eta\%$ of the samples contain only images or only texts, while $(1-\eta)\%$ contain image-text pairs. For the missing-both case, we have $\frac{\eta}{2}\%$ text-only, $\frac{\eta}{2}\%$ image-only, and $(1-\eta)\%$ complete pairs. We follow [24, 25, 26] and set $\eta=70$ throughout our experiments .
We compare our method with (1) baseline multimodal models: ViLT [8] and Visual-BERT [9]; (2) multitask and optimal fusion based method by Ma et al. [3] and (3) prompting based methods: missing-aware prompts (MAP) [24], modality-specific prompts (MSP) [25], multi-step adaptive prompt (MuAP) [26] and method proposed by Kim and Kim [48].
Table 3 shows the results obtained using a consistent missing modality setup during both training and testing. To ensure robust evaluation, we report the mean and standard deviation of three independent runs with different seeds. Our approach outperforms existing baseline methods on both UPMC Food-101 and MM-IMDb datasets across all compared scenarios. These results suggest that the mask tokens can provide good estimates of the missing modality features, which improves missing modality robustness. We also evaluate our model with different modality ratio during training and testing. We discuss the results on Section 3 in the supplementary.
Table 2: Performance (accuracy) comparison on Audio-Video datasets when both modalities are available during training and testing. denotes results generated using existing code.
Method Parameters (M) KS AVE CREMA-D
Total Learnable
G-Blend 22.36 22.36 62.20 65.50 58.70
FiLM 22.35 22.35 63.33 - 60.07
BiGated 22.35 22.35 63.72 - 59.21
OGM-GE 22.35 22.35 65.74 76.90 68.14
QMF 22.36 22.36 65.78  68.16 63.71
PMR 22.36 22.36 - 66.40 61.80
MLA 22.37 22.37 71.35  64.68 79.70
MBT 172.00 172.00 85.00 77.80 -
MMLoRA 172.19 172.19 91.40 96.20 88.60
Ours 172.16 0.17 91.21 96.77 88.84
@@ 4.4.1 Missing modality during both training and testing
Table 3: Performance comparison on MM-IMDb and UPMC Food-101 dataset when modality gets missing during both training and testing. We train and test our model with same modality ratio and report the mean and standard deviation (SD) of three independent runs with different seeds. We report F1-Macro score for MM-IMDb dataset and % accuracy for UPMC Food-101 dataset.
Datasets Missing Train and Test ViLT MAP MSP Visual- Ma Kim MuAP Ours
@@
Rate Image Text (Input) (Attn) BERT et al. et al. (Head) (Cross) (Mean $\pm$ SD)
MM-IMDb 70% 30% 100% 37.61 46.30 44.74 47.45 38.63 46.63 56.03 47.21 46.73 60.21 $\pm$ 0.40
@@
Figure 2: Robustness comparison against missing modalities during test time on MM-IMDb and UPMC Food-101 datasets. Models are trained with 100% image + 100% text and evaluated with 100% image + $x$text. Our approach demonstrates better robustness, particularly under severe modality loss.
Table 4: Performance comparison when a complete modality gets missing during test time. We report % accuracy for all the datasets. Our method shows overall better performance compared to the baseline methods.
Dataset Available Modality Missing Modality FiLM BiGated OGM-GE QMF MLA Ours
@@
Figure 3: Contribution of mask tokens (MT) in model performance when a specific modality gets missing.
4.4.2 Missing modality during testing.
To evaluate the test time missing modality robustness of our approach, we train our models on modality-complete training data and evaluate on modality-incomplete test data. We can divide this into two different scenarios: (i) partial modality missing and (ii) complete modality missing.
Partial modality missing. For this evaluation, we follow the experimental setup similar to [3, 48] where models are trained with 100% image + 100% text data and tested with 100% image + $x$text. We evaluate our model on MM-IMDb and UPMC Food-101 datasets and show the comparison with existing baseline methods proposed by Ma et al. [3], Kim and Kim [48] and ViLT [8] in Figure 2. We see that model performance drops as the amount of available text decreases. Our method demonstrates better robustness compared to other baseline methods. For example, on MM-IMDb dataset, our method achieves 5.8% greater F1-Macro score compared to Kim and Kim [48] when only 10% text available. Consistently, our method outperforms all the baseline methods on all missing scenarios. We see similar trend on UPMC Food-101 dataset where our method shows better performance and thus more robustness when modalities are missing during test time. These methods do not report their performance when different amount of image modality is missing. That is why we could not compare against them in that scenario.
Complete modality missing. We also evaluate our model when a modality is completely missing during test time. For this evaluation, we train our models on complete multimodal data and evaluate when one modality is completely missing during test. As shown in Table 4, our method outperforms existing baseline methods on most of the scenarios. Our method also maintain better robustness compared to other methods. For example, when audio modality gets missing, all the methods except MLA show drastic performance drop on CREMA-D and Kinetics-Sound (KS) datasets. Similar trend is observed when text is missing on UPMC Food-101 dataset. Our method shows less performance drop compared to multimodal performance and also outperforms all the methods. MLA [36] performs better when image is missing on UPMC Food-101 dataset, but shows larger performance drop when text is missing. In contrast, our method maintains a balance in performance when either modality is available. In summary, our method demonstrates overall better performance and robustness in complete modality scenarios.
@@ 4.5 Ablation Studies
Effectiveness of Mask Tokens Figure 3 demonstrates the effectiveness of our proposed Mask Tokens in handling missing modalities on MM-IMDb dataset. The results show consistent performance improvements across all the classes for both text missing and image missing scenarios. Notably, Mask Tokens significantly mitigate performance degradation in modality-sensitive classes. For instance, we see that the performance on “Horror”, “War” and “Animation” classes suffers drastically when text is missing as shown in Figure 3(a). However, by utilizing our proposed Mask Tokens, this performance is improved significantly. This boost in performance is also significant in the “Romance”, “Thriller” and “Action” classes when image is missing as shown in Figure 3(b). By enhancing the per-class performance, Mask Tokens improves the overall model performance when modalities get missing.
Evaluating cross-modal feature estimation. We also evaluate the quality of cross-modal feature estimation of our proposed mask tokens. To do this, we calculate the mean squared error (MSE) between the original features and the features estimated by the mask tokens, and show the per-class statistics on MM-IMDb dataset in Figure 4. Here, A → B indicates that the mask tokens of modality A are used to estimate the features of modality B. As illustrated in the figure, the MSE between the original and estimated features remains consistently low, averaging around 0.04 across all classes. This indicates that Mask Tokens effectively learn to estimate features of the missing modality using information from the available modality. Consequently, the missing modality can be accurately reconstructed, enhancing performance in various missing modality scenarios. The low MSE values highlight the robustness of our approach in handling modality-incomplete data and demonstrate the reliability of our method in estimating missing features.
Analysis of modality dependence. We observe that the impact of missing modalities varies significantly across classes. Certain classes show strong dependence on a specific modality. In this experiment, we train our model on modality-complete data and evaluate it on both modality-complete and modality-incomplete test samples. We present per-class performance on MM-IMDb dataset in Figure 5. Our analysis reveals heterogeneous impacts of missing modalities across different classes. For instance, the “Drama” and “Thriller” classes maintain consistent performance regardless of missing modalities. In contrast, the “Biography” and “Documentary” classes experience significant decline in performance when image is available and text is missing. Interestingly, certain classes like “Animation” show slightly higher performance with single-modality inputs, particularly in image-only scenario. Similar patterns emerge in “Sport” and “Music” categories, where text-only samples performs slightly better than complete modality. This may be due to the simple sum fusion strategy we use, which may not effectively capture complex relationships between modalities. These findings indicate that the effect of missing modalities is not uniform across classes and that some classes are more dependent on certain modalities. A similar pattern is observed in UPMC Food-101 dataset which we discuss in Section 4 in the supplementary.
Figure 4: Per class mean squared error between the real and estimated tokens on MM-IMDb dataset.
5 Discussion and Future Directions
Advantages over existing methods. Despite having some similarities with recent methods, such as MMLoRA [37] and prompt-tuning [24, 25, 26], our method differs from them in both architectural design and training strategy. MMLoRA follows a two-step fine-tuning process. Unimodal models are first fine-tuned on unimodal data, and then LoRA layers are added and adapted to multimodal data. This approach assumes that unimodal fine-tuning is required before multimodal adaptation. In contrast, we demonstrate that jointly fine-tuning unimodal models on multimodal data is sufficient to achieve better or comparable performance across different modalities and multimodal tasks. Additionally, MMLoRA does not address the problem of missing modalities. Our method, on the other hand, introduces Mask Tokens, which effectively handle different missing modality scenarios. Our approach also differs significantly from prompt-tuning based approaches [24, 25, 26]. Prompt-based methods improve missing modality performance by incorporating learnable prompts across multiple layers of the model. They require learning and storing a large number of prompts, typically 16 prompts [24, 25] per modality per layer, for different modality combinations. In contrast, our method uses a single mask token to estimate the missing modality feature from the available modality, thereby reducing the complexity. By fine-tuning the encoders with LoRA, we can effectively estimate the missing modality without the need for learning or storing a large number of prompts. Our approach outperforms prompt based approaches as shown in Table 3.

### v2

@@ 3.4 Gating module and fusion
where $\lambda$ is a hyperparameter that we set to $\lambda=0.20$ in all our experiments.
Our model can handle missing modalities during both training and inference. During training on complete modality data, we apply random modality dropout, where we either use both modalities or randomly drop one in each iteration. This way, the model learns to make correct predictions in missing modality scenarios using the CMPTs from the available modality. When training data contains missing modalities, we learn the CMPTs by optimizing the alignment loss using only the samples with complete modalities, as shown in (4). For missing modalities, we employ placeholders (e.g., an empty string for text and zeros for images, audio, and video) following standard practices [22, 39], and use the CMPT from the available modality as an approximation to make predictions, as described in (5). To enhance training efficiency, we leverage LoRA [15] with a rank of 1 to fine-tune the encoders in a parameter-efficient manner. LoRA layers are inserted after the query, key, value, and output layers in every attention blocks within the encoders. Our experiments demonstrate that LoRA with rank 1 is sufficient to achieve better performance compared to existing state-of-the-art methods, as discussed in Section 4.
4 Experiments and Results
@@
Table 1: Performance comparison when modality gets missing during both training and inference. We train and evaluate our model with same modality ratio following [22, 16, 20] and report the mean and standard deviation (SD) of three independent runs with different seeds.
Datasets % availability ViLT MAP [22] MSP Visual- Ma et al. Kim et al. MuAP [6] CMPT (Ours)
@@
Image Text [21] Input Attn. [16] BERT [23] [29] [20] Head Cross Mean $\pm$ SD
MM-IMDb (F1-Macro) 30% 100% 37.61 46.30 44.74 47.45 38.63 46.63 56.03 47.21 46.73 60.21 $\pm$ 0.40
@@
Table 2: Performance comparison when a complete modality gets missing during inference. We report % accuracy for all the datasets. Our method shows overall better performance compared to recent state-of-the-art methods.
Datasets Available Modality Missing Modality FiLM BiGated OGM-GE QMF MLA CMPT
[37] [18] [36] [54] [55] (Ours)
@@
4.1 Datasets
We evaluate our approach on five popular multimodal datasets across different tasks. A brief description of each dataset is provided here, with further details in Section S2 in the supplementary.
4.2 Implementation details
Audio-Video datasets. We use AST model [13] pretrained on AudioSet dataset [12] as audio encoder and ViT-B [8] from CLIP model [38] as video encoder. Raw audio is converted to a 128-bin mel spectrogram at 16 kHz, with a maximum length of 1024 frames. For video, we sample 3 frames randomly while training and uniformly while testing. Other pre-processing steps are similar to [55, 36, 9].
Hyperparameter settings. We use Python 3.8.19 and PyTorch 2.2.2 for training and evaluating our models. All the models are trained using two NVIDIA RTX 2080Ti GPUs. For vision-language datasets, we set the learning rate to $10^{-3}$ and train the models for 10 epochs with a batch size of 8. For audio-video datasets, the learning rate is set to $5\times 10^{-5}$ and models are trained for 100 epochs with a batch size of 4. We use AdamW [26] optimizer with $\epsilon=10^{-8}$ and weight decay = 0.02. Cross entropy loss is applied while training the models. We utilize a polynomial learning rate scheduler with power=0.9 and set the first 5 epoch as warm-up. The learning rate is set to 0.1 times the original learning rate during warm-up epochs. We set LoRA [15] rank = 1 and insert them after query, key, value and output layers of each transformer block. Further details with all the hyperparameters can be found in Section S3 in the supplementary.
For each task and dataset, we compare against state-of-the-art published work. For this reason, the set of baselines varies across experiments based on the availability of published results. In each table, the best and second best results are bold and underlined, respectively.
Table 3: Performance comparison with both modalities available during training and inference. We report % accuracy for UPMC Food-101 and F1-Macro/F1-Micro score for MM-IMDb dataset. ‘-’ indicate results that are not reported in corresponding papers.
Methods Parameters (M) UPMC MM-IMDb
Total Learnable Food-101
ViLT [21] 112.63 112.63 92.00 55.30 / 64.70
MBT [33] 196.00 196.00 93.56 59.60 / 64.81
MMBT [19] 170.00 170.00 94.10 60.80 / 66.10
MMLoRA [9] 196.10 196.10 93.70 61.70 / 67.20
PMF [24] 198.43 2.54 91.51 58.77 / 64.51
PMF-L [24] 643.95 4.44 91.68 61.66 / 66.72
MLA [55] 218.26 218.26 93.33 - / -
CMPT (Ours) 195.51 0.27 94.47 63.58 / 69.23
4.3 Performance when modalities are missing
Existing literature considers missing modalities at both training and inference or only at inference. Our method achieves state-of-the-art performance in both scenarios, as detailed in the next two sections.
4.3.1 Missing modality during training and inference
In the first setup, a modality is missing at a fixed rate during both training and inference, as studied in [22, 16, 20]. We follow the experiment setup in [22, 16, 20] and compare our method against seven state-of-the-art approaches on the MM-IMDb and UPMC Food-101 datasets. As shown in Table 1, our method consistently outperforms all baselines across various missing modality configurations. Specifically, compared to the most recent approach in Kim et al. [20], we achieve a 4.18%–9.40% improvement in F1-macro score on MM-IMDb dataset and a 0.16%–1.56% improvement in accuracy on UPMC Food-101 dataset across different missing modality scenarios. These results also suggest that the CMPTs can provide good proxy for the missing class tokens, which significanctly improves the model robustness. To ensure a reliable evaluation, we run our method with three different random seeds and report the mean and standard deviation (SD) for various missing scenarios. The low SD values indicate the stability and reliability of our method. We also assessed the performance of the models with different distributions of modality ratios at training and testing, for which the results are presented in Section S4 of the supplementary material.
4.3.2 Missing modality during inference only
In the second setup, all modalities are available during training, but one modality is completely missing during inference, as examined in [55, 39]. Following the setup in [55], we evaluate our model when all modalities are available during training, but one modality is completely missing during inference. As shown in Table 2, our method outperforms existing baselines in most cases. For instance, when the audio modality is missing, all methods except MLA [55] experience a significant performance drop on the CREMA-D and Kinetics-Sound (KS) datasets. On the CREMA-D dataset, our method achieves 7.93% and 11.30% improvements over the most recent MLA approach, when video and audio are missing, respectively. Similarly, on the KS dataset, our method outperforms MLA with 13.60% and 34.74% improvements when video and audio are missing, respectively. While MLA performs better when the image modality is missing on the UPMC Food-101 dataset, it suffers a larger performance drop when text is missing. In contrast, our method maintains a balanced performance across different missing modalities. Furthermore, across all datasets, our method achieves the best performance when all modalities are available and demonstrates superior robustness compared to other approaches. These results highlight the effectiveness of our method in handling scenarios where a modality is entirely absent during inference.
Table 4: Performance (accuracy) comparison with both modalities available during training and inference on Audio-Video datasets. denotes results generated using existing code.
Method Parameters (M) KS AVE CREMA-D
Total Learnable
G-Blend [45] 22.36 22.36 62.20 65.50 58.70
FiLM [37] 22.35 22.35 63.33 - 60.07
BiGated [18] 22.35 22.35 63.72 - 59.21
OGM-GE [36] 22.35 22.35 65.74 76.90 68.14
QMF [54] 22.36 22.36 65.78  68.16 63.71
PMR [11] 22.36 22.36 - 66.40 61.80
MLA [55] 22.37 22.37 71.35  64.68 79.70
MBT [19] 172.00 172.00 85.00 77.80 -
MMLoRA [9] 172.19 172.19 91.40 96.20 88.60
CMPT (Ours) 172.16 0.17 91.21 96.77 88.84
(a) Dataset: MM-IMDb
(b) Dataset: UPMC Food-101
Figure 3: Generalization to varying missing rates during inference. Models are trained with 100% image + 100% text and evaluated with 100% image + $x$text following [29, 20]. Our approach demonstrates better generalization, particularly under severe modality loss.
4.4 Performance when all modalities are available
A method designed to handle missing modalities should not compromise performance when all modalities are present. In this subsection, we demonstrate that our approach also achieves comparable or superior performance to state-of-the-art models across five datasets when all modalities are available. These results highlights that incorporating CMPTs does not introduce any drawback when all modalities remain available.
To evaluate this, we train and test our method when all modalities are available during training and inference. Results in Table 3 demonstrate that our proposed method achieves state-of-the-art performance on both the UPMC Food-101 and MM-IMDb datasets. Our method surpasses the most recent MLA approach [55] by 1.14% on UPMC Food-101. Our method achieves 1.88% and 2.03% improvement in F1-macro and F1-micro scores, respectively, over second-best performing MMLoRA [9] on MM-IMDb.
We present the results for audio-video datasets in Table 4, where we compare our method against nine state-of-the-art methods. Our approach consistently achieves comparable or superior performance across all three datasets, while using the fewest number of learnable parameters.
4.5 Generalization to different missing rates
In real-world scenarios, the missing rates for modalities during inference can be unpredictable. Any method designed to handle missing modalities needs to generalize effectively across varying levels of missing modality rates during inference. To evaluate this capability, we trained our models on complete multimodal data and assessed their performance under varying levels of missing modalities during inference. Specifically, we follow the experiment setup similar to [29, 20] where models are trained with 100% image + 100% text data and evaluated with 100% image + $x$text.
As illustrated in Figure 3, our method consistently outperforms all baselines across different missing scenarios, demonstrating superior robustness. We observe a decline in performance of all methods as the amount of available text decreases, but our approach maintains stronger resilience compared to other methods. For instance, on the MM-IMDb dataset, our method achieves a 5.8% higher F1-Macro score than the most recent method by Kim et al. [20] and a 25.6% improvement over ViLT [21] when only 10% of text is available. A similar trend is observed on the UPMC Food-101 dataset, where our method consistently yields superior performance, confirming its robustness to missing modalities at inference.
Note that published work in [29, 20, 21] did not report results for different missing rates of images, which prevents us from reporting a direct comparison in that setting. Nevertheless, our results highlight the effectiveness of CMPTs in handling different missing modality scenarios, making it a more generalizable solution for real-world applications.
(a) Available: Text, Missing: Image.
(b) Available: Image, Missing: Text.
Figure 4: Effectiveness of CMPT in handling missing modalities on MM-IMDb dataset. CMPT significantly improves performance when modalities are severely missing. All models are trained on 100% image + 100% text and tested with varying levels of missing modalities.
@@ 4.5 Generalization to different missing rates
Figure 5: CMPT improves performance across most of the classes when modalities are missing. However, it fails to accurately predict certain classes (e.g., short/documentary) in the absence of text, likely due to the image modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
@@ 4.6 Ablation studies
4.6.1 Effectiveness of CMPT
To evaluate the effectiveness of CMPTs in handling missing modalities, we compare our method against two baselines: (i) a standard model trained without modality dropout or CMPT (baseline) and (ii) a model trained with modality dropout, where modalities are randomly removed during training to improve robustness to missing data.
Figure 4 presents the main results for MM-IMDb dataset, where the baseline model experiences a sharp performance drop as the proportion of missing modalities increases. While modality dropout improves robustness, it fails to fully recover the lost performance. Incorporating CMPT further enhances performance, achieving a 32.06% improvement over the baseline and 11.55% over modality dropout when the image modality is entirely missing as shown in Figure 4(a). A similar trend is observed when text is missing as shown in Figure 4(b). Notably, CMPTs become increasingly effective as the amount of missing data increases, demonstrating its strong ability to mitigate performance degradation in the presence of severe modality loss. Similar trend appears in UPMC Food-101 dataset, which we discuss in Sec. S5 in the supplementary material.
4.6.2 Effects of CMPT on different classes
When analyzing the effectiveness of CMPT, a key question is whether the overall performance gains are driven by just a few classes or if CMPT consistently improves performance across most classes. To address this, we conducted a per-class performance analysis with and without CMPT.
As shown in Figure 5, CMPT improves performance across nearly all classes in both text-missing and image-missing scenarios. Notably, it significantly mitigates performance degradation in modality-sensitive classes. For example, when text is missing, classes like Horror, Animation, and War experience severe drops in performance, but CMPT effectively recovers much of the loss (Figure 5(a)). Similarly, when images are missing, CMPT provides substantial gains in Thriller, Action, and Romance (Figure 5(b)). A slight drop is observed in Short and Documentary classes when text is missing, likely due to insufficient information in the image modality alone for those classes. These results demonstrate that CMPT enhances performance across most classes, leading to a strong overall performance improvement. A similar trend is observed on the UPMC Food-101 dataset (further discussed in Section S6 of the supplementary material).
5 Limitations and Future Directions